# Ultraluminous X-ray sources in Globular Clusters

### v1

@@
The results indicate that the apparent scarcity of ULXs in GCs may be related to the older stellar populations of GCs relative to the field. Furthermore, the study identifies a population of ”escaper” ULXs, which originate in GCs but are ejected and emit X-rays outside the cluster. These escapers may significantly contribute to the observed field ULX population.
@@ 1 Introduction
Ultraluminous X-ray sources (ULXs) have been a subject of intense study in the field of high-energy astrophysics for several decades. These enigmatic objects, characterized by X-ray luminosities exceeding $10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$, challenge our understanding of accretion physics and compact object formation (see e.g. Kaaret et al. 2017a, b, for a recent review). With over $1800$ known ULXs identified to date (Walton et al. 2022), they represent a significant population of extreme X-ray emitters in the Universe, yet their fundamental nature remains a topic of ongoing debate.
@@
The diversity of ULXs has been highlighted in recent classification schemes (e.g. Wiktorowicz et al. 2017), which attempt to categorize these objects based on their observational properties and potential underlying physical mechanisms. While the majority of ULXs were initially thought to harbor intermediate-mass black holes (IMBHs), our understanding has evolved considerably with the detection of pulsar accretors (Bachetti et al. 2014).
@@
The importance of studying ULXs, particularly in the context of globular clusters (GC), cannot be overstated. These objects serve as laboratories for extreme accretion physics, potentially shedding light on the formation and evolution of double compact objects (DCOs) - the progenitors of gravitational wave sources. Moreover, ULXs continue to be considered as candidates for elusive IMBHs (Kaaret et al. 2017a, and references therein).
A critical aspect of ULX research involves identifying and characterizing their donor stars. Heida et al. (2016); López et al. (2020) reported that some ULXs had been identified with potential red super-gian (RSG) donors using infrared spectroscopy data. They examined a total of $113$ ULXs using photometric techniques in the infrared band, resulting in the identification of candidate counterparts for $38$ objects. Among these, the nature of $12$ sources was confirmed through spectroscopic analysis, with only five categorized as ULXs likely to have RSG donors.
@@
In this paper, we present a comprehensive study of ULXs in globular clusters, focusing on their stellar-mass manifestations. The dense stellar environment introduces additional complexities and possibilities for exotic binary formations. By examining their properties, distribution, and potential formation mechanisms, we aim to contribute to the broader understanding of ULX physics and their role in compact object evolution.
1.1 Globular Cluster environment
GCs are dense, quasi-spherical collections of stars that orbit the centers of galaxies. These typically old stellar systems, often containing millions of stars within a relatively small volume, are hubs of dynamical interactions. Their evolution is governed by the relaxation process connected with continuous distance gravitational interactions between all objects in the cluster.
@@ 1.1 Globular Cluster environment
This leads to global phenomena like, core collapse, i.e. contraction of the cluster core and gradual increase in central density, and mass segregation, i.e. the tendency of heavier stars to occupy the central regions, while pushing the lighter stars outwards. Close dynamical interactions influence also the stars and binaries individually, influencing the binary parameters (e.g. the separation and eccentricity), or through exchanges (e.g. binary changes one of it’s component into another star), ejections (star or binary becomes unbound with the cluster), and binary formation (single stars became bound and form a binary). Such dynamical encounters in dense star clusters can also lead to the formation of close binary systems, including X-ray binaries (Pooley et al. 2003; Ivanova et al. 2005, 2008). Understanding how the fundamental dynamics influence the ULX formation is essential for interpreting the presence and characteristics of these exotic objects in GC environments.
Figure 1: A schematic representation of the origins and eventual locations of ULX progenitors and ULXs in GCs. Progenitors can either form and remain bound to the cluster, leading to in-cluster ULX emission (”IN-CLUSTER”), or be dynamically ejected and emit as ULXs outside the cluster (”ESCAPER”). Additionally, ULXs originating and emitting in the field (”FIELD”) represent progenitors formed without direct interaction with cluster dynamics.
1.2 GCULX candidates
@@
While the majority of ULXs are detected in star-forming regions of galaxies, there is increasing evidence for their presence in GCs as well. Pioneering observations of GCULXs have been reported in several galaxies, with varied GC environments, spectral behavior, and variability patterns.
@@
Maccarone et al. (2007) discovered the first ULX in a GC (NGC 4472). It exhibits an X-ray luminosity of $4\times 10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$ and rapid variability. Similar detections were reported in NGC 1399 (Shih et al. 2010; Irwin et al. 2010), with more GCULXs in NGC 4472 (Maccarone et al. 2011) and NGC 4649 (Roberts et al. 2012). Some sources identified in GCs exhibited flares just above the ULX-defining limit of $10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$ (Irwin et al. 2016; Sivakoff et al. 2005).
@@
Recent observations have extended the known GCULX population, including seven additional GCULXs associated with M87 GCs (Dage et al. 2020), three in NGC 1316 (Dage et al. 2021), and $10$ new candidates in GCs of massive ($>10^{11.5}M_{\odot}$) early-type galaxies (Thygesen et al. 2023).
@@
In dense stellar systems, compact objects can be readily ejected, which can considerably impact the formation channels of such objects (e.g. Kulkarni et al. 1993). Furthermore, GCs typically exhibit high ages and, consequently, fewer higher-mass stars. These characteristics contribute to the intriguing nature of GCULXs.
@@ 2.1 MOCCA Monte Carlo Code
The MOCCA (MOnte Carlo Cluster evolution Code; Hypki & Giersz 2013) is a state-of-the-art numerical tool that combines the Monte Carlo method for stellar dynamics (Hénon 1971; Stodolkiewicz 1986) with detailed stellar and binary evolution algorithms from the BSE code (Hurley et al. 2000, 2002) with recent updates (level-C from Kamlah et al. 2022, and references therein) and low-N scattering code FEWBODY (Fregeau et al. 2004) to follow close dynamical interactions. This hybrid approach makes MOCCA particularly well-suited for studying the formation and evolution of exotic objects like ULXs in dense stellar environments, offering an optimal balance between computational efficiency and physical accuracy.
Recent improvements to the MOCCA code (Hypki et al. 2022, 2025; Giersz et al. 2024) have significantly enhanced its capabilities for ULX studies. First of all, the mass transfer calculations now incorporate a more sophisticated treatment of super-Eddington accretion, crucial for modeling ULX systems (Shakura & Sunyaev 1973; Wiktorowicz et al. 2015). Additionally, improvements in core radii calculations improved the RLOF calculations. The introduction of detailed evolutionary tracking provide unprecedented insight into the spatial distribution and evolutionary pathways of ULX systems within clusters equivalent to detailed outputs from codes like startrack (Belczynski et al. 2008).
A major advancement relevant to this study is MOCCA’s capability to handle multiple stellar populations (MSPs), a phenomenon now recognized as ubiquitous in globular clusters (e.g. Piotto et al. 2015; Renzini et al. 2015; Gratton et al. 2019; Bastian & Lardo 2018; Milone & Marino 2022). While a comprehensive analysis of MSPs in our models is presented in Hypki et al. (2022, 2025); Giersz et al. (2024), here we specifically investigate their impact on ULX formation and evolution. The code now tracks different stellar populations with distinct chemical compositions, ages, and spatial distributions, allowing us to examine how MSP-specific properties affect compact object formation and ULX characteristics.
All recent changes to the MOCCA code are described in details in Giersz et al. (2024).
@@ 2.2 Simulations
The MOCCA code employs hundreds of parameters to govern the dynamical and stellar evolution of globular clusters. Based on previous studies of compact object populations - including cataclysmic variables (Belloni et al. 2017), ULXs in the field (Wiktorowicz et al. 2019), and double white dwarf systems (Hellström et al. 2024) - we identified key parameters that most significantly influence ULX formation and evolution. Our parameter selection strategy focused on those that control dynamics and initial stellar population properties. This systematic approach allows us to explore the most relevant parameter space while maintaining computational feasibility. In this section we present discussion of the utilized parameters.
@@
For all simulations with dynamics enabled, we tracked the evolution of escaped systems to analyze the properties of their populations and potential ULXs that can form after their progenitors are ejected from the cluster. These escapers evolve as isolated binaries, similar to field binary evolution (e.g., Fragos et al. 2015; Wiktorowicz et al. 2019; Zuo et al. 2021). Systems can escape the cluster through multiple mechanisms: (1) dynamical interactions, particularly strong few-body encounters, (2) natal kicks from supernovae explosions, including both direct and Blaauw kicks, (3) gradual two-body relaxation (Spitzer & Hart 1971), and (4) tidal stripping of outer cluster regions by the host galaxy’s gravitational field (Gnedin & Ostriker 1997).
@@
The galactocentric distance ($\mathrm{R}_{\mathrm{gc}}$) for each simulation was scaled to maintain an initial tidal radius of $\mathrm{R}_{\mathrm{tid}}\approx 43\mathrel{\mathrm{pc}}$ across all models. This normalization of the tidal radius facilitates meaningful comparisons between simulations, particularly when analyzing the degree of tidal filling and cluster structural parameters. The tidal radius follows the relation $\mathrm{R}_{\mathrm{tid}}\propto\mathrm{R}_{\mathrm{gc}}^{2/3}\mathrm{M}_{\mathrm{tot}}^{1/3}$ (e.g. Webb et al. 2013), where $\mathrm{M}_{\mathrm{tot}}$ is the cluster mass. This scaling ensures that clusters experience comparable relative tidal forces despite different masses and orbital parameters. We note that while the initial tidal radii are identical, the clusters’ subsequent evolution may lead to different filling factors111filling factor is R_tid/R_h, which is initially $\sim 43$ for non-tidally filling simulations and $\sim 3.6$ for tidally filling ones. depending on their internal dynamics and mass-loss history.
Table 1: Model parameters and initial properties
npop1-fb10-TF 1 10% $2.01$ $11.90$ $0.69$ $5.85$ $89.10$ $2.64$ $10,888.00$
npop1-fb10-nTF 1 10% $2.01$ $1.00$ $0.69$ $0.49$ $89.10$ $5.87$ $265.22$
npop1-fb95-TF 1 95% $1.96$ $11.93$ $0.72$ $5.81$ $93.93$ $2.61$ $6,386.70$
npop1-fb95-nTF 1 95% $1.96$ $1.00$ $0.72$ $0.49$ $93.93$ $5.84$ $155.06$
npop2-cpop05-fb95-nTF 2 0.05 95% $1.96$ $1.00$ $0.69$ $0.01$ $13.03$ $9.59$ $93.85$
npop2-cpop2-fb10-nTF 2 0.2 10% $2.01$ $1.00$ $0.66$ $0.04$ $12.84$ $7.77$ $174.58$
npop2-cpop2-fb95-nTF 2 0.2 95% $1.96$ $1.00$ $0.69$ $0.04$ $13.12$ $7.77$ $101.01$
222 List of simulations with differentiating parameters and some initial properties (at $t=0$). Label identifies particular simulation and will be used throughout the paper. TF/nTF in the label means initially tidally filling or non-tidally filling. NF, when present, means that new features were included in this model (see text for detailes); $\mathrm{N}_{\mathrm{pop}}$ number of stellar generations; $\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$ ratio between the $\mathrm{R}_{\mathrm{h}}$ of the second and first stellar generation; $\mathrm{f}_{\mathrm{bin}}$ binary fraction; $\mathrm{R}_{\mathrm{gc}}$ galactocentric radius [$\mathrel{\mathrm{kpc}}$]; $\mathrm{R}_{\mathrm{h}}$ half-mass radius for the first population [$\mathrel{\mathrm{pc}}$]; $\mathrm{M}_{\mathrm{tot}}$ total stellar mass [$\mathrel{\mathrm{M}_{\odot}}$]; $\mathrm{R}_{\mathrm{c}}$ core radius [$\mathrel{\mathrm{pc}}$]; $\mathrm{M}_{\mathrm{c}}$ core mass [$\mathrel{\mathrm{M}_{\odot}}$]; $\rho_{\mathrm{c}}$ central density [$\mathrel{\mathrm{M}_{\odot}}/\mathrel{\mathrm{pc}}^{3}$]; $\mathrm{t}_{\mathrm{h,rel}}$ Spitzer half-mass relaxation time [$\mathrel{\mathrm{M}_{\odot}}$]. All simulations include the escapers information (see text). For all these simulations the counterpart field population (with dynamics turned off) was calculated (nodyn in the label; not shown in the table) with the same parameters.
@@
We performed a comprehensive set of simulations exploring different initial conditions and physical parameters. Table 1 summarizes our simulation grid, where each model is identified by a unique label used throughout this paper. The simulations vary in the number of stellar populations ($\mathrm{N}_{\mathrm{pop}}$), with varied spatial distribution of different generations through the concentration parameter ($\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$), defined as the ratio between the half-mass radii of subsequent generation relative to the first generation. The initial binary fraction ($\mathrm{f}_{\mathrm{bin}}$) was either $10\%$, or $95\%$, which affected also other parameters (see below). We explored both tidally filling (TF) and non-tidally filling (nTF) configurations ($\mathrm{Tidal}$). Some simulations incorporate the new features (NF) described below, particularly relevant for multiple population scenarios. For each simulation, we provide key structural parameters at $t=0\mathrel{\mathrm{Myr}}$, including the galactocentric radius ($\mathrm{R}_{\mathrm{gc}}$), half-mass radius ($\mathrm{R}_{\mathrm{h}}$), total stellar mass ($\mathrm{M}_{\mathrm{tot}}$), core radius ($\mathrm{R}_{\mathrm{c}}$; according to Casertano & Hut 1985), core mass ($\mathrm{M}_{\mathrm{c}}$, i.e. mass inside $\mathrm{R}_{\mathrm{c}}$), and central density ($\rho_{\mathrm{c}}$; according to Casertano & Hut 1985). Additionally, we computed corresponding field populations with dynamics turned off (denoted by nodyn in the model labels) to serve as control cases, though these are not shown separately in the table.
@@
To better compare simulations with different binary fractions, we made several adjustments to the parameter space. Since a lower binary fraction results in lower total stellar mass (as binary systems typically have higher masses than single stars), we adjusted the initial number of objects to have similar half-mass relaxation time for both models: simulations with $\mathrm{f}_{\mathrm{bin}}=95\%$ used $n=600,000$ objects, while those with $\mathrm{f}_{\mathrm{bin}}=10\%$ used $n=1,063,635$ objects.We also modified the initial distribution of semi-major axes. For simulations with $\mathrm{f}_{\mathrm{bin}}=95\%$, we used the modified version of the Kroupa (1995) period distribution (see Belloni et al. 2017) for stars with $M<5\mathrel{\mathrm{M}_{\odot}}$, and the distributions from Sana et al. (2012); Oh et al. (2015) for stars with $M>5\mathrel{\mathrm{M}_{\odot}}$. For simulations with $\mathrm{f}_{\mathrm{bin}}=10\%$, we employed a log-uniform distribution for $M<5\mathrel{\mathrm{M}_{\odot}}$ and the Sana et al. (2012) period distribution for $M>5\mathrel{\mathrm{M}_{\odot}}$. Additionally, for the low binary fraction simulations, we disabled eigenevolution (Belloni et al. 2017). The separation was limited to the maximal value of $100\mathrel{\mathrm{AU}}$
@@
In a part of our simulations (marked with NF in label), we introduced several modifications to investigate the influence of multiple stellar populations (see Giersz et al. 2024, for details). Specifically, we implemented a delay time ($t_{\rm delay}$) of $100\mathrel{\mathrm{Myr}}$ for the second population, during which these stars act as gas particles without undergoing stellar evolution, relaxation, or dynamical interactions. The gas accumulation for the second population begins at $t_{\rm delay\_fraction}=0.5$, meaning halfway through the delay period, coinciding with the asymptotic giant branch (AGB) wind contribution phase. To accommodate this, we activated AGB wind accumulation for the second population formation. The cluster’s orbital evolution was modified by implementing an artificial orbit change at $t=1000\mathrel{\mathrm{Myr}}$, where the galactocentric distance increases by a factor of 2 (see Giersz et al. 2024, for details). Finally, we employed single-population scaling, where the scaling parameters are applied exclusively to the first population, while the second population’s velocities are independently normalized. First and second populations are separately in the virial equilibrium.

### v3

@@
The results indicate that the apparent scarcity of ULXs in GCs may be related to the older stellar populations of GCs relative to the field. Furthermore, the study identifies a population of ”escaper” ULXs, which originate in GCs but are ejected and emit X-rays outside the cluster. Our simulations reveal that these escapers constitute about one-seventh of the total ULX population. However, for neutron star accretors specifically, escapers are twice as common as their in-cluster counterparts. Notably, only $4\%$ of in-cluster ULXs contain neutron star accretors. These escapers may significantly contribute to the observed field ULX population.
@@ 1 Introduction
Ultraluminous X-ray sources (ULXs) have been a subject of intense study in the field of high-energy astrophysics for several decades. These enigmatic objects, characterized by X-ray luminosities exceeding $10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$, challenge our understanding of accretion physics and compact object formation (see e.g., Kaaret et al., 2017; King et al., 2023, for a recent review). With over $1800$ known ULXs identified to date (Walton et al., 2022), they represent a significant population of extreme X-ray emitters in the Universe, yet their fundamental nature remains a topic of ongoing debate.
@@
The diversity of ULXs has been highlighted in recent classification schemes (e.g., Wiktorowicz et al., 2017), which attempt to categorize these objects based on their observational properties and potential underlying physical mechanisms. While the majority of ULXs were initially thought to harbor intermediate-mass black holes (IMBHs), our understanding has evolved considerably with the detection of pulsar accretors (Bachetti et al., 2014).
@@
The importance of studying ULXs, particularly in the context of globular clusters (GC), cannot be overstated. ULXs serve as laboratories for extreme accretion physics, potentially shedding light on the formation and evolution of double compact objects - the progenitors of gravitational wave sources. Moreover, ULXs continue to be considered as candidates for elusive IMBHs with masses in the range of $10^{2}$ to up to $10^{5}\mathrel{\mathrm{M}_{\odot}}$ (Kaaret et al., 2017, and references therein). The existence of IMBHs may provide crucial insights into the formation pathways of supermassive BHs observed at the centers of most galaxies. These intermediate-mass objects could also explain the presence of supermassive BHs observed at high redshift (Greene et al., 2020; Askar et al., 2023).
@@
In this paper, we present a comprehensive study of ULXs in GCs, focusing on their stellar-mass manifestations. The dense stellar environment introduces additional complexities and possibilities for exotic binary formations. By examining their properties, distribution, and potential formation mechanisms, we aim to contribute to the broader understanding of ULX physics and their role in compact object evolution.
Dense, quasi-spherical collections of stars known as GCs orbit the centers of galaxies. These typically old stellar systems, often containing millions of stars within a relatively small volume, are hubs of dynamical interactions. The masses of Milky Way GCs range from a few $10^{3}$ to a few $10^{6}\ \mathrel{\mathrm{M}_{\odot}}$, with half-mass radii typically between 1.5 pc and a few tens of pc, and ages spanning from 8 to 13.5 Gyr (Harris, 1996; Salaris & Weiss, 2002; VandenBerg et al., 2013; Baumgardt & Hilker, 2018). Detailed properties of specific Galactic GCs can be found in the catalogs of Harris (1996, updated 2010) [^1] https://physics.mcmaster.ca/~harris/mwgc.dat and Baumgardt & Hilker (2018) [^2] https://people.smp.uq.edu.au/HolgerBaumgardt/globular/parameter.html.
@@
The evolution of GCs is governed by the relaxation process connected with continuous distant gravitational interactions between all objects in the cluster. This leads to global phenomena like core collapse, that is contraction of the cluster core and gradual increase in central density, and mass segregation, that is the tendency of heavier stars to occupy the central regions while pushing lighter stars outwards. Close dynamical interactions also influence stars and binaries individually, affecting binary parameters (e.g., separation and eccentricity), or through exchanges (e.g., binary changes one of it’s component into another star), ejections (star or binary becomes unbound with the cluster), and binary formation (single stars became bound and form a binary).
Such dynamical encounters in dense star clusters can lead to the formation of close binary systems, including X-ray binaries (Pooley et al., 2003; Ivanova et al., 2005, 2008; Kremer et al., 2018, 2019), cataclysmic variables (Oh et al., 2024; Belloni et al., 2019), and gravitational wave progenitors (e.g., Benacquista & Downing, 2013; Mandel & Broekgaarden, 2022, and references therein). Recent studies have emphasized the importance of GC dynamics in shaping exotic stellar populations (Di Carlo et al., 2021; Oh et al., 2024), highlighting the role of these environments in producing and evolving compact object binaries.
@@
While the majority of ULXs are detected in star-forming regions of galaxies, there is increasing evidence for their presence in GCs as well. These GCULXs are found exclusively in extragalactic GCs, all of which are unresolved at distances of ¿16 Mpc, with no ULXs detected in Milky Way GCs to date. Pioneering observations of GCULXs have been reported in several galaxies, with varied GC environments, spectral behavior, and variability patterns.
@@
Maccarone et al. (2007) discovered the first ULX in a GC (NGC 4472). It exhibits an X-ray luminosity of $4\times 10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$ and rapid variability. Similar detections were reported in NGC 1399 (Shih et al., 2010; Irwin et al., 2010), with more GCULXs in NGC 4472 (Maccarone et al., 2011) and NGC 4649 (Roberts et al., 2012). Some sources identified in GCs exhibited flares just above the ULX-defining limit of $10^{39}\mathrel{\mathrm{erg}/\mathrm{s}}$ (Irwin et al., 2016; Sivakoff et al., 2005).
@@
Recent observations have extended the known GCULX population, including seven additional GCULXs associated with M87 GCs (Dage et al., 2020), three in NGC 1316 (Dage et al., 2021), $10$ new candidates in GCs of massive ($>10^{11.5}M_{\odot}$) early-type galaxies (Thygesen et al., 2023), and two more in NGC 4261 (Nair et al., 2023). The current population of approximately 30 known GCULXs, identified through systematic searches of thousands of GC systems in elliptical galaxies, represents a remarkably rare phenomenon compared to the $\sim 1800$ ULXs identified in other galactic environments (Walton et al., 2022).
@@
In dense stellar systems, compact objects can be readily ejected, which can considerably impact the formation channels of such objects (e.g., Kulkarni et al., 1993). Furthermore, GCs typically exhibit high ages and, consequently, fewer higher-mass stars. Given their presence in older stellar populations, GCULXs are likely powered by low-mass X-ray binaries, in contrast to field ULXs which are typically associated with massive donor stars but often lack optical counterparts. Given their presence in older stellar populations, GCULXs are likely powered by low-mass X-ray binaries, in contrast to field ULXs which are expected to be also associated with low-mass as well as massive donor stars, but often lack optical counterparts. These characteristics contribute to the intriguing nature of GCULXs.
@@ 2.1 MOCCA Monte Carlo code
The MOCCA (MOnte Carlo Cluster evolution Code; Hypki & Giersz, 2013) is a state-of-the-art numerical tool that combines the Monte Carlo method for stellar dynamics (Hénon, 1971; Stodolkiewicz, 1986) with detailed stellar and binary evolution algorithms from the Binary Star Evolution code (BSE; Hurley et al., 2000, 2002) with recent updates (level-C from Kamlah et al., 2022, and references therein) and low-N scattering code FEWBODY (Fregeau et al., 2004) to follow close dynamical interactions. This hybrid approach makes MOCCA particularly well-suited for studying the formation and evolution of exotic objects like ULXs in dense stellar environments, offering an optimal balance between computational efficiency and physical accuracy.
Recent improvements to the MOCCA code (Hypki et al., 2022, 2025; Giersz et al., 2024) have significantly enhanced its capabilities for ULX studies. First of all, the mass transfer calculations now incorporate a more sophisticated treatment of super-Eddington accretion, crucial for modeling ULX systems (Shakura & Sunyaev, 1973; Wiktorowicz et al., 2015). Additionally, improvements in core radii calculations improved the Roche-lobe overflow calculations. The introduction of detailed evolutionary tracking provide unprecedented insight into the spatial distribution and evolutionary pathways of ULX systems within clusters equivalent to detailed outputs from codes like startrack (Belczynski et al., 2008).
A major advancement relevant to this study is MOCCA’s capability to handle multiple stellar populations, a phenomenon now recognized as ubiquitous in GCs (e.g., Piotto et al., 2015; Renzini et al., 2015; Gratton et al., 2019; Bastian & Lardo, 2018; Milone & Marino, 2022). While a comprehensive analysis of multiple stellar populations in our models is presented in Hypki et al. (2022, 2025); Giersz et al. (2024), here we specifically investigate their impact on ULX formation and evolution. The code now tracks different stellar populations with distinct chemical compositions, ages, and spatial distributions, allowing us to examine how properties specific to multiple stellar populations affect compact object formation and ULX characteristics. For detailed analysis of simulational results on multiple stellar populations and all recent changes to the MOCCA code see Giersz et al. (2024, and references therein).
The X-ray luminosity ($\mathrm{L}_{\mathrm{X}}$) of accretion disks is calculated using the Shakura-Sunyaev model (Shakura & Sunyaev, 1973), accounting for both sub-Eddington and super-Eddington accretion regimes. The Eddington mass accretion rate is defined as $\dot{\mathrm{M}}_{\mathrm{Edd}}=2.2\times 10^{-8}\mathrm{M}_{\mathrm{acc}}\eta_{0.1}^{-1}\;[\mathrel{\mathrel{\mathrm{M}_{\odot}}\;\mathrm{yr}^{-1}}]$, where $\mathrm{M}_{\mathrm{acc}}$ is the accretor mass and $\eta\;(=10\eta_{0.1})$ is the accretion efficiency. We assumed $\eta_{0.1}=1$. The Eddington luminosity is then $\mathrm{L}_{\mathrm{Edd}}\approx 5.66\times 10^{45}\;\dot{\mathrm{M}}_{\mathrm{Edd}}\;[\text{erg s}^{-1}]$. The X-ray luminosity is calculated as:
$\mathrm{L}_{\mathrm{X}}=\cases{\mathrm{L}_{\mathrm{Edd}}}(\dot{\mathrm{M}}/\dot{\mathrm{M}}_{\mathrm{Edd}}),&\text{for }\dot{\mathrm{M}}\leq\dot{\mathrm{M}}_{\mathrm{Edd}}\\ \mathrm{L}_{\mathrm{Edd}}[1+\ln(\dot{\mathrm{M}}/\dot{\mathrm{M}}_{\mathrm{Edd}})],\text{for }\dot{\mathrm{M}}>\dot{\mathrm{M}}_{\mathrm{Edd}}{}$ (1)
where $\dot{\mathrm{M}}$ is the mass accretion rate. See Lasota & King (2023) for recent discussion of this formulation.
@@ 2.2 Simulations
Figure 1: Schematic representation of the origins and eventual locations of ULX progenitors and ULXs in GCs. Progenitors can either form and remain bound to the cluster, leading to in-cluster ULX emission (”IN-CLUSTER”), or be dynamically ejected and emit as ULXs outside the cluster (”ESCAPER”). Additionally, ULXs originating and emitting in the field (”FIELD”) represent progenitors formed without direct interaction with cluster dynamics.
The MOCCA code employs hundreds of parameters to govern the dynamical and stellar evolution of GCs. Based on previous studies of compact object populations - including cataclysmic variables (Belloni et al., 2017), ULXs in the field (Wiktorowicz et al., 2019), and double white dwarf systems (Hellström et al., 2024) - we identified key parameters that most significantly influence ULX formation and evolution. Our parameter selection strategy focused on those that control dynamics and initial stellar population properties. This systematic approach allows us to explore the most relevant parameter space while maintaining computational feasibility. In this section we present discussion of the utilized parameters.
@@
For all simulations with dynamics enabled, we tracked the evolution of escaped systems to analyze the properties of their populations and potential ULXs that can form after their progenitors are ejected from the cluster. These escapers evolve as isolated binaries, similar to field binary evolution (e.g., Fragos et al., 2015; Wiktorowicz et al., 2019; Zuo et al., 2021). Systems can escape the cluster through multiple mechanisms: (1) dynamical interactions, particularly strong few-body encounters, (2) natal kicks from supernovae explosions, including both direct and Blaauw kicks, (3) gradual two-body relaxation (Spitzer & Hart, 1971), and (4) tidal stripping of outer cluster regions by the host galaxy’s gravitational field (Gnedin & Ostriker, 1997).
@@
The galactocentric distance ($\mathrm{R}_{\mathrm{gc}}$) for each simulation was scaled to maintain an initial tidal radius of $\mathrm{R}_{\mathrm{tid}}\approx 43\mathrel{\mathrm{pc}}$ across all models. This normalization of the tidal radius facilitates meaningful comparisons between simulations, particularly when analyzing the degree of tidal filling and cluster structural parameters. The tidal radius follows the relation $\mathrm{R}_{\mathrm{tid}}\propto\mathrm{R}_{\mathrm{gc}}^{2/3}\mathrm{M}_{\mathrm{tot}}^{1/3}$ (e.g., Webb et al., 2013), where $\mathrm{M}_{\mathrm{tot}}$ is the cluster mass. This scaling ensures that clusters experience comparable relative tidal forces despite different masses and orbital parameters. We note that while the initial tidal radii are identical, the clusters’ subsequent evolution may lead to different filling factors333filling factor is $\mathrm{R}_{\mathrm{tid}}$/$\mathrm{R}_{\mathrm{h}}$, which is initially $\sim 43$ for non-tidally filling simulations and $\sim 3.6$ for tidally filling ones. depending on their internal dynamics and mass-loss history.
@@
We performed a comprehensive set of simulations exploring different initial conditions and physical parameters. Table 7 summarizes our simulation grid, where each model is identified by a unique label used throughout this paper. The simulations vary in the number of stellar populations ($\mathrm{N}_{\mathrm{pop}}$), with varied spatial distribution of different generations through the concentration parameter ($\mathrm{R}_{\mathrm{h,2}}/\mathrm{R}_{\mathrm{h,1}}$), defined as the ratio between the half-mass radii of subsequent generation relative to the first generation. The initial binary fraction ($\mathrm{f}_{\mathrm{bin}}$) was either $10\%$, or $95\%$, which affected also other parameters (see below). We explored both tidally filling (TF) and non-tidally filling (nTF) configurations ($\mathrm{Tidal}$). Some simulations incorporate the new features (NF) described below, particularly relevant for multiple population scenarios. For each simulation, we provide key structural parameters at $t=0\mathrel{\mathrm{Myr}}$, including the galactocentric radius ($\mathrm{R}_{\mathrm{gc}}$), half-mass radius ($\mathrm{R}_{\mathrm{h}}$), total stellar mass ($\mathrm{M}_{\mathrm{tot}}$), core radius ($\mathrm{R}_{\mathrm{c}}$; according to Casertano & Hut, 1985), core mass ($\mathrm{M}_{\mathrm{c}}$, i.e., mass inside $\mathrm{R}_{\mathrm{c}}$), and central density ($\rho_{\mathrm{c}}$; according to Casertano & Hut, 1985). Additionally, we computed corresponding field populations with dynamics turned off (denoted by nodyn in the model labels) to serve as control cases, though these are not shown separately in the table.
@@
To better compare simulations with different binary fractions, we made several adjustments to the parameter space. Since a lower binary fraction results in lower total stellar mass (as binary systems typically have higher masses than single stars), we adjusted the initial number of objects to have a similar half-mass relaxation time for both models: simulations with $\mathrm{f}_{\mathrm{bin}}=95\%$ used $n=600,000$ objects, while those with $\mathrm{f}_{\mathrm{bin}}=10\%$ used $n=1,063,635$ objects. We also modified the initial distribution of semi-major axes. For simulations with $\mathrm{f}_{\mathrm{bin}}=95\%$, we used the modified version of the Kroupa (1995) period distribution (see Belloni et al., 2017) for stars with $M<5\mathrel{\mathrm{M}_{\odot}}$, and the distributions from Sana et al. (2012); Oh et al. (2015) for stars with $M>5\mathrel{\mathrm{M}_{\odot}}$. For simulations with $\mathrm{f}_{\mathrm{bin}}=10\%$, we employed a log-uniform distribution for $M<5\mathrel{\mathrm{M}_{\odot}}$ and the Sana et al. (2012) period distribution for $M>5\mathrel{\mathrm{M}_{\odot}}$. Additionally, for the low binary fraction simulations, we disabled eigenevolution (Belloni et al., 2017). The separation was limited to the maximal value of $100\mathrel{\mathrm{AU}}$
@@
In a part of our simulations (marked with NF in label), we introduced NF to investigate the influence of multiple stellar populations (see Giersz et al., 2024, for details). Specifically, we implemented a delay time ($t_{\rm delay}$) of $100\mathrel{\mathrm{Myr}}$ for the second population, during which these stars act as gas particles without undergoing stellar evolution, relaxation, or dynamical interactions. The gas accumulation for the second population begins at $t_{\rm delay\_fraction}=0.5$, meaning halfway through the delay period, coinciding with the asymptotic giant branch (AGB) wind contribution phase. To accommodate this, we activated AGB wind accumulation for the second population formation.
The cluster’s orbital evolution was modified by implementing an artificial orbit change at $t=1000\mathrel{\mathrm{Myr}}$, where the galactocentric distance increases by a factor of 2 (see Giersz et al., 2024, for details). Finally, we employed single-population scaling, where the scaling parameters are applied exclusively to the first population, while the second population’s velocities are independently normalized. First and second populations are separately in the virial equilibrium.